sir,in sine rule a by sinA=b by sinB=c by sinC=2R ,a,b,c are side of triangle&A,B,C are angle of a triangle then what is R?

Grade:12th Pass

2 Answers

Chetan Mandayam Nayakar
12 years ago

S=√(s(s-a)(s-b)(s-c)),s=(a+b+c)/2

R=abc/4S=radius of circumcircle

KAMAL RAWAT
12 years ago

R=abc/4s

where s=a+b+c/3